what is probability that a king will be selected on the first draw from the deck and another king on the second draw(assuming that the first king was not replaced)
Answer by stanbon(75887)

what is probability that a king will be selected on the first draw from the deck and another king on the second draw(assuming that the first king was not replaced)
----
# of ways to succeed: 4C2 = 6
--
# of possible outcomes: 52C2 = 1326
---
P(king and then king) = 6/1326 = 3/663 = 1/221
--------------------
Cheers,
Stan H.